# Can MLLMs Guide Me Home? A Benchmark Study on Fine-Grained Visual Reasoning from Transit Maps

### v2

@@ D.2 Limitations and Future Work
While ReasonMap provides a carefully curated benchmark for evaluating fine-grained visual reasoning with high-resolution transit maps, we acknowledge that it represents only one type of structured visual diagram. As such, caution should be taken when generalizing observations to other domains that involve different types of visual content or reasoning styles. Additionally, although efforts were made to ensure diversity across cities and languages, the current version may not fully capture all geographic or linguistic variations. Future iterations could further expand coverage and explore additional forms of reasoning to enhance generality.
D.3 Broader Impact
@@
Appendix E License and Consent Information
E.1 Ethical Considerations
All experiments are conducted on ReasonMap, which is built using publicly available transit maps collected in compliance with relevant licenses and usage terms. The maps are selected to ensure geographic diversity and legal validity. Upon code release, we provide the source of each map for further reference. ReasonMap is intended solely for academic research on fine-grained visual understanding and spatial reasoning in MLLMs. It does not redistribute any copyrighted map images. All annotations are based on public information, contain no personal data, and are created under academic oversight. The benchmark is not intended for safety-critical use. We take care to ensure fairness, legal compliance, and responsible data handling.
E.2 Public Implementation

### v3

@@ E.1 Limitations and Future Work
While ReasonMap provides a carefully curated benchmark for evaluating fine-grained visual reasoning with high-resolution transit maps, we acknowledge that it represents only one type of structured visual diagram. As such, caution should be taken when generalizing observations to other domains that involve different types of visual content or reasoning styles. Additionally, although efforts were made to ensure diversity across cities and languages, the current version may not fully capture all geographic or linguistic variations. Future iterations could further expand coverage and explore additional forms of reasoning [17] to enhance generality.
Furthermore, we note that GeoGuessr-style localization tasks [45, 25, 30] are compelling, as they emphasize detailed visual understanding of natural scenes and signage. We plan to pair transit maps with street view imagery to support cross-view reasoning and localization within ReasonMap, thereby expanding beyond static map inputs. In parallel, we will explore agent-based training and evaluation that moves from single-turn prediction to iterative planning with feedback, including reward designs for correctness, calibration, and format [99]. Finally, we will extend toward embodied settings [28] where agents perceive and act in interactive environments, enabling assessment of instruction following, route planning, and navigation under real-world constraints. Together, these directions broaden the benchmark from fine-grained visual reasoning to context-aware spatial intelligence and practical decision making.
Our ReasonMap can further evaluate the efficient models from multiple efficiency strategies [101, 55, 56, 62, 63, 14]. Additionally, more fields [97, 98, 2, 72, 32, 38, 39, 8, 34, 66, 41, 91, 93, 94] require corresponding reasoning-centered benchmarks for proper evaluation.
E.2 Broader Impact
@@
Appendix F Further Statement
F.1 Public Implementation
@@
F.2 Large Language Model Usage Statement
We used a large language model (LLM) solely for surface-level editing of the manuscript (e.g., rephrasing for clarity and concision, grammar/style polishing, and minor fixes). The LLM did not generate technical content, ideas, algorithms, proofs, code, experiments, figures, or tables; the authors conducted all research design, implementation, data processing, and analyses. The model did not produce or select citations; any suggestions were independently verified and replaced with primary sources. Interactions were limited to de-identified text snippets of the manuscript, and no non-public data, code, or unreleased results were uploaded. All LLM outputs were manually reviewed and edited by the authors. This usage does not affect reproducibility: every reported number is reproducible from our released code and configurations.
F.3 Ethics Statement
All experiments are conducted on ReasonMap, which is built using publicly available transit maps collected in compliance with relevant licenses and usage terms. The maps are selected to ensure geographic diversity and legal validity. Upon code release, we provide the source of each map for further reference. ReasonMap is intended solely for academic research on fine-grained visual understanding and spatial reasoning in MLLMs. It does not redistribute any copyrighted map images. All annotations are based on public information, contain no personal data, and are created under academic oversight. The benchmark is not intended for safety-critical use. We take care to ensure fairness, legal compliance, and responsible data handling. Additionally, we will use the MIT License for code release on GitHub and the Apache License 2.0 for ReasonMap release on HuggingFace.